OpenStudy (anonymous):

I need help converting 375mL to cm^3

OpenStudy (aaronq):

\(\sf mL=cm^3\) 100 cm = 1 m so, \(\sf 375~cm^3*(\dfrac{1~m}{100~cm})^3=0.000375~m^3\)

OpenStudy (anonymous):

@aaronq It's supposed to end up with _____ cm^3

OpenStudy (aaronq):

sorry, you wrote "3m^3" first, so i thought the 3 was a typo. it's the same thing, \(\sf 1 ~mL=1~cm^3\)

OpenStudy (anonymous):

Yeah, sorry about that. Thanks!

OpenStudy (aaronq):

no problem!